OpenStudy (anonymous):

solve the equation please 3\2b+8=20

mathslover (mathslover):

I hope that this is the question : \[\huge{\frac{3}{2b}+8=20}\] is it ?

OpenStudy (anonymous):

it is but the b is beside the 3/2 haha

mathslover (mathslover):

u mean : \[\huge{\frac{3b}{2}+8=20}\]

OpenStudy (anonymous):

\[\frac{3}{2}b + 8 = 20\]\[\frac32b = 12\]\[b = \frac{12}{1}\times\frac{2}{3}\]\[b = ?\]This works because dividing by a fraction is the same as multiplying by it's reciprocal.

OpenStudy (anonymous):

do i cross mulitiply to figure out what b =? Calcmathlete?